A ladder is 1 foot longer than its vertical height along the side of the house. The distance of the base of the
ladder from the house is 7 feet less than its vertical height from the ground. How far up the house does the
ladder reach?

Answer by jorel1380(3719)

x2+(x-7)2=(x+1)2
x2+x2-14x+49=x2+2x+1
x2-16x+48=0
(x-12)(x-4)=0
x=12 or 4
Throwing out the smaller answer, we get the height along the wall to be 12 ft., the distance of the ladder from the wall to be 5 ft., and the height of the ladder to be 13 ft...